# Evaluating the Relationship Between News Source Sharing and Political Beliefs

### v2

@@ 1 Introduction
In this work, correspondence analysis is employed not to infer users’ ideology but to quantify their media preferences based on the news articles they share. We compare these preferences with their political ideology previously inferred through a machine learning model developed and validated in Zhou et al. (2021) based only on the hashtags supporting one of the two coalitions that contested the elections that year [zhou2021]. Specifically, we use users’ connections to Argentinean news articles to deduce preferences for media outlets within a latent space, achieved by conducting correspondence analysis of the user-media matrix. The comparison of both metrics allows us to measure how the political ideology of the users is related to the preference for the media they share. Additionally, we delve into the emergence of community structures within the retweet network (excluding retweets containing links to news articles) to discern whether user interactions are driven by political leaning or media outlet preferences.
@@ 2.1 Argentinian Context
Beyond the back and forth of the political context in Argentina, there is still a lack of quantitative analysis of the relation between political ideology an media preferences, specially during the social media age.
@@ 3.1 Data description
In this research, we employed a pre-existing Twitter dataset [zhou2021] comprising tweets collected between March 1, 2019, and August 27, 2019. As it is explained in [zhou2021], this dataset was acquired by utilizing the following keywords linked to participants in the 2019 Argentinean primary election: Alberto AND Fernández, alferdez, CFK, CFKArgentina, Kirchner, mauriciomacri, Macri, Pichetto, MiguelPichetto, Lavagna. Exclusively, tweets written in Spanish were selected.
First, we filtered for all types of tweets, including tweets, retweets, and quotes, that exclusively contained external URLs linking to Argentinian news outlets, based on the ABYZ News Links Guide [abyz]. Following these initial steps, we obtained a dataset encompassing the activity of 123,180 users, who collectively generated 1,039,281 tweets, sharing 66,982 unique news articles from 17 Argentinian media outlets.
@@ 3.2.1 Correspondence Analysis
Following previous studies [barbera2015tweeting, flamino2023political] that propose a methodology for inferring user coordinates in a latent space of social media based on correspondence analysis [greenacre2010correspondence], we begin by establishing a bipartite network denoted as G = (U, V, E), where U represents the set of users, V denotes the news outlets, and E stands for the edges in the graph. The corresponding adjacency matrix associated with this network is denoted as Y. The element $y_{ij}$ represents the number of times user $i$, with $i\in U$ shares a news from a media outlet $j$, with $j\in V$. The main difference between our implementation and the seminal proposal [barbera2015tweeting] is that the bipartite network here is that the connections between users and media comes from the content of the users’ tweets and not from explicit network connections (e.g., following-followers relations or retweets).
@@ 4.1 Media Sharing Index
As described above, we construct the bipartite adjacency matrix $Y$, where each element $y_{ij}$ represents the number of news articles from media outlet $j$ shared by user $i$. We apply correspondence analysis to this matrix to calculate the Media Sharing Index (MSI), as discussed in section 3.2.1. For simplicity, we focus on the primary 12 media outlets, thereby reducing the dataset to 59,874 unique news articles (approximately 88% of the total unique news in the dataset) and 120,626 users (about 97% of the users in the original dataset). These users originate from a total of 1,015,380 tweets containing links to one of these 12 main outlets, which constitutes about 98% of the original tweet volume.
@@ 4.2 Media sharing and political preferences
The asymmetric behavior observed between ideological groups can be explained by noticing that media outlets with MSI $>0$, such as Clarín, La Nación, and Infobae, are also those that dominate the Argentine media landscape [comscore] (see section 2.1). This observation reveals that, despite political preferences, center-left users can not overlook these outlets given their important role in the political discussion on social media.
@@ 4.3 Retweet user networks
The association found between the community label and the ideological position of users supports the reproduction of the results discussed in section 4.2 at a community level: the center-right community exhibits media sharing behavior favoring center-right media outlets, as depicted in the blue histogram of the MSI. Conversely, the center-left community displays a less biased MSI distribution, indicating that this community shares content from both center-left media outlets and those biased towards the opposite ideological spectrum. As mentioned in section 4.2, this diversity seems to arise because CR media outlets represent mainstream media, and the CL group cannot afford to overlook them when engaging in political discourse.
@@ 5 Discussion and Conclusions
The strength of the Media Sharing Index lies in its ability to quantify news sharing behavior among diverse media outlet groups. Coupled with the inference of users’ political leaning derived from the machine learning model outlined in [zhou2021], we can explore the correlations between these dimensions and address questions such as: Can the bias of the shared media outlets be used as a proxy of user’s ideology? Our analysis unveils non-trivial findings: while the CR group predominantly shares news from media aligned with their own beliefs, the CL group exhibits greater diversity in its media sources. This diversity arises from the fact that CR media outlets represent mainstream media, and the CL group cannot overlook them when engaging in political discourse.
When we analyze interactions among users, we see the emergence of two large groups. Each one of these groups can be identified with a given political stance (center-right and center-left) but not necessarily with their media sharing preferences. While in the center-right community, political leanings align with a preference for sharing right-biased media, this alignment does not occur in the center-left community, where the distribution of MSI points indicates that individuals share news from media outlets across the political spectrum. This reinforces the fact that the center-left (CL) community cannot ignore center-right (CR) media outlets, especially considering that, in this Argentinian case, they represent the most consumed media outlets in the country.

### v3

@@ 1 Introduction
In this work, correspondence analysis is employed not to infer users’ ideology but to quantify their media preferences based on the news articles they share. We compare these preferences with their political ideology previously inferred through a machine learning model developed and validated in Zhou et al. (2021) based only on the hashtags supporting one of the two coalitions that contested the elections that year [zhou2021]. Specifically, we use users’ connections to Argentinean news articles to deduce preferences for media outlets within a latent space, achieved by conducting correspondence analysis of the user-media matrix. The comparison of both metrics allows us to measure how the political ideology of the users is related to the preference for the media they share. Additionally, we delve into the emergence of community structures within the retweet network (excluding retweets containing links to news articles) to discern whether user interactions are driven by political leaning or media outlet preferences. This work aims to contribute to the quantitative analysis of the relationship between political ideology and media preferences, continuing the line of previous studies such as [calvo2023winning] and [aruguete2023news].
@@ 3.1 Data description
In this research, we employed a pre-existing Twitter dataset [zhou2021] comprising tweets collected between March 1, 2019, and August 27, 2019. The details of this dataset can be found in Appendix A. From this dataset, we filtered for all types of tweets, including tweets, retweets, and quotes, that exclusively contained external URLs linking to Argentinian media outlets listed in the ABYZ News Links Guide [abyz]. From this list we selected 17 media outlets by ranking the outlets in descending order by the number of times they were shared on Twitter and selecting those with broad recognition and influence not only on this platform but also across Argentine media as a whole, as reported in sources such as [digitalnewsreport]. Given this filter, we first obtained a dataset encompassing the activity of 123,180 users, who collectively generated 1,039,281 tweets, sharing 66,982 unique news articles.
@@ 3.2.1 Correspondence Analysis
Following previous studies [barbera2015tweeting, flamino2023political] that propose a methodology for inferring user coordinates in a latent space of social media based on correspondence analysis [greenacre2010correspondence], we begin by establishing a bipartite network denoted as G = (U, V, E), where U represents the set of users, V denotes the news outlets, and E stands for the edges in the graph. The corresponding adjacency matrix associated with this network is denoted as Y. The element $y_{ij}$ represents the number of times user $i$, with $i\in U$ shares a news from a media outlet $j$, with $j\in V$. The main difference between our implementation and the seminal proposal [barbera2015tweeting] is that the bipartite network here is based on the content of users’ tweets, rather than explicit network connections (e.g., following-followers relations or retweets).
@@ 4.1 Media Sharing Index
As described above, we construct the bipartite adjacency matrix $Y$, where each element $y_{ij}$ represents the number of news articles from media outlet $j$ shared by user $i$. We apply correspondence analysis to this matrix to calculate the Media Sharing Index (MSI), as discussed in section 3.2.1. For simplicity, we focus on the primary 12 media outlets, excluding media outlets that are shared by only a few users and where the majority of the shared articles come from the outlets themselves. This reduces the dataset to 59,874 unique news articles (approximately 88% of the total unique news in the dataset) and 120,626 users (about 97% of the users in the original dataset). These users originate from a total of 1,015,380 tweets containing links to one of these 12 main outlets, which constitutes about 98% of the original tweet volume.
@@ 4.2 Media sharing and political preferences
A possible interpretation of the asymmetric behavior observed between ideological groups is that media outlets with MSI $>0$, such as Clarín, La Nación, and Infobae, are also predominant players in the Argentine media landscape [comscore] (see section 2.1). These outlets’ extensive reach and influence may contribute to the observed sharing patterns among center-left users at the aggregate level.
@@ 4.3 Retweet user networks
The association found between the community label and the ideological position of users supports the reproduction of the results discussed in section 4.2 at a community level: the center-right community exhibits media sharing behavior favoring center-right media outlets, as depicted in the blue histogram of the MSI. Conversely, the center-left community displays a less biased MSI distribution, indicating that this community shares content from both center-left media outlets and those biased towards the opposite ideological spectrum. As mentioned in section 4.2, the observed diversity in sharing patterns among the CL group may be linked to the prominence of CR media outlets in the overall media ecosystem.
@@ 5 Discussion and Conclusions
The strength of the Media Sharing Index lies in its ability to quantify news sharing behavior among diverse media outlet groups. Coupled with the inference of users’ political leaning derived from the machine learning model outlined in [zhou2021], we can explore the correlations between these dimensions and address questions such as: Can the bias of the shared media outlets be used as a proxy of user’s ideology?
Our analysis reveals distinct patterns: while the CR group predominantly shares news from media that align with their ideological stance, the CL group exhibits a broader range of media sources in their sharing behavior. This observed diversity could likely be influenced by the prominent role of CR media outlets within the broader media ecosystem in Argentina. Accordingly, this ideological asymmetry in news preferences and exposure has been also widely documented in other contexts. For instance, in the US, studies have examined the media ecosystem [benkler2018network, jamieson2008echo] and the preference for news exposure in social media. In [hmielowski2020asymmetry], the authors analyze the reinforcing effects of liberal and conservative media on political beliefs during the 2016 US election, finding that conservative beliefs contribute more to a conservative media echo chamber than liberal beliefs do to a liberal one. Similarly, [gonzalez2023asymmetric] explores ideological segregation in political news on Facebook during the 2020 US election, revealing that conservatives consume a disproportionately large amount of news, including most misinformation, while liberal sources are less prevalent. Additionally, findings regarding the sharing of misinformation and fake news indicate that, during the 2016 US election, conservative Facebook users were more likely to share pro-Trump fake news than their liberal or moderate counterparts [guess2019less]. Research conducted in countries outside the US also suggests that ideology has a stronger influence on right-wing users compared to left-wing users when it comes to news-sharing behavior [aruguete2019news].
@@
Appendix A Original dataset
As mentioned in 3.1, we employed a pre-existing Twitter dataset [zhou2021] comprising tweets collected between March 1, 2019, and August 27, 2019. This dataset was obtained through the Twitter Academic API (at the time of acquiring the data, this API allowed full access to the complete history of the social media platform) by searching for the following queries: Pichetto, Espert, victoria, Bregman, argentina, Voto, Kirchner, Propuesta Republicana, Moreno, Peron, Sola, kirchnerismo, Massa, UCR, PASO, Alternativa Federal, Justicialista, PRO, Movimiento Socialismo, Rossi, #ArgentinaVota, Union Civica Radical, fuerza, Votar, Vidal, Alfonsin, alferdez, kicillof, Scioli, vamos, Frente Izquierda, Donda, miguelpichetto, CFK, Vote, kicillofok, Alberto Fernandez, Urtubey, elecciones, sergiomassa, ganamos, Peronista, CFKArgentina, peronismo, #EleccionesArgentina, PJ, Lavagna, Partido Justicialista, mauriciomacri, Consenso 19, Cristina, Lousteau, Cambiemos, macrismo, mariuvidal, Capitanich, apoyo, Macri. Exclusively, tweets written in Spanish were selected.
The full dataset analyzed comprised around 170 million tweets and about 3.5 million users. Although the dataset is massive, given the keywords used to obtain it, the dataset may include users with a certain bias towards tweeting about politics and, therefore, must be considered with caution when evaluating its representativeness of the Argentinian population. On the other hand, Zhou et al. [zhou2021] reported an average classification per day of 114,653 tweets as CR supporters and 96,576 as CL supporters (MP and FF, respectively, in [zhou2021]), which indicates a balanced classification.
All of these numbers are drastically reduced when considering only those tweets with URLs linking to Argentinian outlets, as stated in section 3.1.